# Understanding the difficulty of low-precision post-training quantization of large language models

### v1

@@ 4.2 Scaling with LLM size and numerical precision
Next, we questioned how the generalization abilities of quantized models produced by RTN, QAFT and GPTQ, scaled as a function of model size and numerical precision. In order to gain a comprehensive understanding of the matter, we aggregated all models in a single tradeoff graph of test $\mathrm{NLL}$ loss plotted against total transformer weight sizes in gigabytes (Figure 2).
@@
As shown in Figure 2, the Pareto front was dominantly occupied by networks quantized through QAFT. In the top panel for the GPT-2 model family, the Pareto front mainly consisted of solutions by QAFT with int6 and int8. At a same model size, a GPT-2 model quantized to int6 and int8 outperformed both a full-precision counterpart with fewer parameters, and one quantized to lower precision with more parameters. We observed a similar pattern with models from the OPT and Llama 2 families, as shown in the lower panel of Figure 2. Here, int4, int6 and int8 models produced by QAFT dominated the Pareto front.
@@
In both panels of Figure 2, we highlighted the difference in test $\mathrm{NLL}$ between QAFT and GPTQ solutions, by the colored, transparent and vertical strips. Evidently, models quantized by QAFT was better generalizing at lower precision, such as int2, int3 and int4. This observation was consistent with the misalignment between the global $\mathrm{NLL}$ loss and the layer-wise local $\mathrm{MSE}$ losses was more prominent at low numerical precision.
4.3 Explanation of the misalignment from a loss landscape perspective
@@
Finally, we asked why minimization of local layer-wise $\mathrm{MSE}$ losses did not align with minimizing the global $\mathrm{NLL}$ loss, especially in case of quantization to low numerical precision. To answer this question, we resorted to mapping of the global $\mathrm{NLL}$ loss landscape in the $D$-dimensional weight space in the neighborhood of the pretrained weights ${\bm{w}}\in{\mathbb{R}}^{D}$.
We presented an example in Figure 3 of opt-125m, a model small enough for numerous loss evaluations.
@@
First, we measured $\mathrm{NLL}$ loss along a number of random directions emanating from ${\bm{w}}$ (the thin, light gray lines in Figure 3). It was evident from the radial mapping that ${\bm{w}}$ sat near the bottom of an attractive basin. The near-convergence neighborhood looked like quadratic, and the loss profiles along randomly sampled directions were rather similar to each other. This is consistent with prior analysis of the loss Hessian spectra showing the presence of a large bulk subspace and a small number of outliers (Sagun et al., 2017, 2018; Ghorbani et al., 2019); thusly, a random linear combination of Hessian eigenvectors would, with high probability, stay within the bulk subspace, making independent radial loss profile look very similar to each other. Beyond the near-convergence locality, notably, the radial loss landscape quickly deviated from a good second-order approximation and plateaued at a high level, demarcating an $\ell_{2}$ neighborhood size $R({\bm{w}})$ that defined the attractive basin.
@@
Next, we charted the loss landscape of the quantized weights via RTN, QAFT and GPTQ, and of linear interpolation segments between them (colored circles and line segments in the Figure 3). We made the following observations.
@@
  • ${\bm{w}}_{\textnormal{RTN}}$ (blue circle), the quantized weight nearest to ${\bm{w}}$, was within the attractive basin for int8 and int6, clearly outside of it for int3 and int2, and around the border for int4.
@@
  • ${\bm{w}}_{\textnormal{GPTQ}}$ (red circle), was significantly more distant from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, but the direction from ${\bm{w}}$ leading to ${\bm{w}}_{\textnormal{GPTQ}}$ (red segment) was much less steep than any typical directions. The loss values at ${\bm{w}}_{\textnormal{GPTQ}}$ were low for int8 and int6, and high for int4, int3 and int2, not much different from those at ${\bm{w}}_{\textnormal{RTN}}$ for the same precision.
@@
  • ${\bm{w}}_{\textnormal{QAFT}}$ (green circle), was not materially more distant from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, but at a significantly lower loss level.
  • Notably, all quantized weights more distant from ${\bm{w}}$ than the radius of the attractive basin, did not seem to be simply connected with ${\bm{w}}$. By simply connected, we meant that the loss profile along the linear segment connecting the two is monotonic without a ridge in the middle.
Based on these observations of the analysed loss landscape, we were able to explain the experimental findings reported in previous sections. The minimization of layer-wise local $\mathrm{MSE}$ by GPTQ, as expected, was successful in picking a much less steeply rising direction emanating from ${\bm{w}}$, along which, if the quantized ${\bm{w}}_{\textnormal{GPTQ}}$ was not distant, stayed at a lower loss level, as in the cases for int8 and int6; but if the quantized ${\bm{w}}_{\textnormal{GPTQ}}$ was substantially distant from ${\bm{w}}$, as in the cases for int3 and int2, the initial shallow rise near ${\bm{w}}$ no longer mattered, resulting in very high loss values at the eventual ${\bm{w}}_{\textnormal{GPTQ}}$ instead. In contrast, even at localities distant from ${\bm{w}}$, QAFT still made material progress from ${\bm{w}}_{\textnormal{RTN}}$, leading to ${\bm{w}}_{\textnormal{QAFT}}$ with significantly lower loss, albeit no longer simply connected to ${\bm{w}}$, likely in another distinct attractive basin, reminiscent of similar observations in training of sparse neural networks (Evci et al., 2020).
@@
Our results suggested that, the relative relationship between the sizes of the attractive basin near convergence $R({\bm{w}})$, and of the weight perturbation due to quantization $\lVert\Delta{\bm{w}}\rVert=\lVert{\bm{w}}_{\textnormal{RTN}}-{\bm{w}}\rVert$, seemed to predict how much minimizing local losses and minimizing the global loss would align with each other. When the latter was substantially larger than the former, misalignment invariably occurred.
5 Discussion
@@
In this work, we conducted a systematic study contrasting and comparing the effectiveness of post-training quantization method GPTQ and quantization-aware fine-tuning (QAFT); the former minimized layer-wise local quantization errors whereas the latter minimized the global training loss. We discovered that, with our experimental setting under the same low training data constraint ($128$ training examples), GPTQ was not as effective as QAFT in nearly all cases. Additional experiments showed that, in certain cases, just a few QAFT iterations could overtake GPTQ in producing a better generalizing LLM (see Appendix B).
@@
We found, rather surprisingly, that in the cases where GPTQ fared significantly worse than QAFT, the minimization of the local $\mathrm{MSE}$ and the global $\mathrm{NLL}$ losses did not necessarily align with each other. We further elucidated, through loss landscape analysis, that such misalignment was a characteristic behavior when the weight perturbation due to quantization was significantly larger than the radius of the attractive basin at the pretraining convergence, a frequent situation consequent from very low precision weight quantization.
@@
Our study demonstrated an apparent lack of correlation between the generalization capabilities of a quantized neural network and the local quantization errors in its component layers, the latter of which often used as a metric to evaluate the quality of quantization schemes. We call for caution on generalization of the utility of local error minimization-based post-training quantization techniques in practice, and provided a new perspective toward understanding the difficulty and identification of the circumstances under which applying these methods were effective.
Appendix A Optimal learning rate for QAFT
While performing hyperparameter grid search on initial learning rate over the set $\{10^{-6},10^{-5},10^{-4},10^{-3}\}$, we discovered that the optimal choices strongly correlated with the model size and the quantization precision, which is supported by Fig 4. For a particular model, we observe that QAFT at higher precision such as int8, int6 and int4 favored low learning rates, whereas QAFT at lower precision such as int3 and int2 preferred high learning rates. For a specific format, larger models in the same model family preferred lower learning rates while smaller models in the model family prefers higher learning rates.
Figure 4: Left: validation $\mathrm{NLL}$ loss of gpt2-large after fine-tuning. Dashed lines represent the loss after RTN. Right: best learning rates for GPT-2 family.
Appendix B Effect of number of QAFT iterations
Although we used 8 epochs ($1024$ iterations) to achieve the most optimal performance, we discovered that QAFT for $1$ epoch is sufficient to outperform GPTQ, as shown in Figure 5.
Figure 5: $\mathrm{NLL}$ on test data after each epoch of fine-tuning. Epoch 00 represents the $\mathrm{NLL}$ of RTN, which is the starting point for fine-tuning. The horizontal red line marks the $\mathrm{NLL}$ after GPTQ.
Appendix C Optimal GPTQ dampening factor
In GPTQ, dampening factor is multiplied with the average diagonal value in Hessian matrix $H$ and added to the diagonal entries of $H$ to achieve better numerical stability [Frantar et al., 2022a]. In our experiments, we did a grid search on dampening factor for each layer with the objective of minimizing layer-wise $\mathrm{MSE}$. We visualized the best dampening factor for each quantized layer but failed to find any meaningful patterns.
Appendix D Additional plots for GPTQ vs. QAFT
Appendix E Additional plots for loss landscape
Figure 6: Left: validation loss landscape for opt-350m. Right: validation loss landscape for opt-1.3b. Figure legend consistent with Figure 3.

### v2

@@ IV-B Scaling with LLM size and numerical precision
Next, we examined how the generalization abilities of quantized models produced by RTN, QAFT and GPTQ, scaled with model size and numerical precision. To capture this, we plotted test $\mathrm{NLL}$ loss against total transformer weight size (in gigabytes) for all models on a single trade-off graph (Figure 2).
@@
As shown in Figure 2, QAFT consistently dominated the Pareto front across model families. For GPT-2 models, QAFT quantized solutions at int6 and int8 occupied the Pareto front, outperforming both smaller full-precision models and larger models quantized with lower precisions. Similar trends were observed for the OPT and Llama 2 families, where QAFT quantized models at int4, int6, and int8 dominated.
@@
Colored vertical strips in Figure 2 highlight test $\mathrm{NLL}$ differences between QAFT and GPTQ. Models quantized by QAFT outperforms GPTQ more significantly at lower precisions, such as int2, int3 and int4, aligning with the earlier observation that misalignment between global $\mathrm{NLL}$ and local $\mathrm{MSE}$ losses is more pronounced at low numerical precision.
IV-C Explanation of the misalignment from a loss landscape perspective
@@
Finally, we investigated why minimization of local layer-wise $\mathrm{MSE}$ losses did not align with minimizing the global $\mathrm{NLL}$ loss, especially at low numerical precision. To address this, we analyzed the global $\mathrm{NLL}$ loss landscape in the $D$-dimensional weight space around the pretrained weights ${\bm{w}}\in{\mathbb{R}}^{D}$.
@@
We measured $\mathrm{NLL}$ loss along a number of random directions emanating from ${\bm{w}}$ (the thin, light gray lines in Figure 3). The radial mapping revealed that ${\bm{w}}$ sat near the bottom of an attractive basin. Within this near-convergence region, the loss landscape appeared quadratic, with similar profiles across various random directions. This aligns with prior analysis of the loss Hessian spectra showing a dominant bulk subspace with a few outliers  [34, 35, 11]. Hence, a random linear combination of Hessian eigenvectors would likely stay within the bulk subspace, resulting in similar radial profiles. Beyond the near-convergence locality, the loss landscape deviated from a quadratic approximation and plateaued at a high level, defining the attractive basin’s radius $R({\bm{w}})$.
@@
Next, we charted the loss landscape of the quantized weights via RTN, QAFT and GPTQ, and of linear interpolation segments between them (colored circles and line segments in Figure 3). Key observations include:
@@
  • ${\bm{w}}_{\textnormal{RTN}}$ (blue circle): The closest quantized weight to ${\bm{w}}$, It was within the attractive basin for int8 and int6, near the border for int4, and outside it for int3 and int2.
@@
  • ${\bm{w}}_{\textnormal{GPTQ}}$ (red circle): Further from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, but along a flatter loss direction (red segment). The loss values at ${\bm{w}}_{\textnormal{GPTQ}}$ were low for int8 and int6, but high for int4, int3 and int2, similar to ${\bm{w}}_{\textnormal{RTN}}$ at the same precision.
@@
  • ${\bm{w}}_{\textnormal{QAFT}}$ (green circle): Slightly further from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, yet achieving significantly lower loss level.
  • Connectivity: Quantized weights beyond the basin’s radius were not simply connected to ${\bm{w}}$. Linear interpolation between ${\bm{w}}$ and these weights showed non-monotonic loss profiles with ridges in the middle.
Based on these observations, we are able to explain experimental findings from previous sections. Minimizing layer-wise local $\mathrm{MSE}$ by GPTQ effectively identified less steeply rising directions from ${\bm{w}}$, resulting in lower loss levels when ${\bm{w}}_{\textnormal{GPTQ}}$ remained within the attractive basin (e.g. int8 and int6). However, for larger quantization-induced perturbations, the shallow rise near ${\bm{w}}$ is insufficient, resulting in high loss values at ${\bm{w}}_{\textnormal{GPTQ}}$ (e.g. int3 and int2). In contrast, QAFT consistently follow less steep directions from ${\bm{w}}_{\textnormal{RTN}}$, even when ${\bm{w}}_{\textnormal{QAFT}}$ was distance from ${\bm{w}}$. This led to significantly lower loss levels, albeit in a separate attractive basin, similar to patterns observed in sparse networks [7].
@@
Our results suggested that the alignment between local and global loss minimization depends on the relationship between the attractive basin size $R({\bm{w}})$ and the quantization-induced weight perturbation $\lVert\Delta{\bm{w}}\rVert=\lVert{\bm{w}}_{\textnormal{RTN}}-{\bm{w}}\rVert$. When $\lVert\Delta{\bm{w}}\rVert$ is substantially greater than $R({\bm{w}})$, misalignment occurs.
IV-D Generalization on different datasets
To assess the generalizability of our findings, we repeated a subset of experiments on C4 [6] and LAMBADA [30] datasets. In both cases, we observed consistent misalignment between the minimization of global and local loss (Figure 4), as well as similar trends in the loss landscape (Figure 5).
Figure 4: Misalignment between minimization of the global NLL loss (by QAFT) and minimization of the local layer-wise MSE losses (by GPTQ) on different datasets. Follows the same conventions as Figure 1.
Figure 5: Loss landscape analysis of quantized model weights on different datasets. Plots showing results for opt-125m quantized in int4. Follows the same conventions as Figure 3.
V Discussion
@@
In this work, we systematically compared the effectiveness of post-training quantization GPTQ and quantization-aware fine-tuning (QAFT), where GPTQ minimized layer-wise local quantization errors and QAFT minimized the global training loss. Under the same low training data constraint ($128$ training examples), QAFT consistently outperformed GPTQ.
@@
Surprisingly, we observed that GPTQ’s local $\mathrm{MSE}$ minimization and QAFT’s global $\mathrm{NLL}$ minimization often misalign, especially pronounced at very low precision quantization. Through loss landscape analysis, we elucidated that such misalignment was prominent when the weight perturbation due to quantization was significantly larger than the radius of the attractive basin at the pretraining convergence.
@@
Our findings reveal a lack of correlation between a quantized network’s generalization ability and its local quantization errors, challenging the common reliance on local error minimization metrics for evaluating quantization schemes. We urge caution in generalizing the utility of local-error-based post-training quantization and provided a new perspective for understanding the difficulty and identification of conditions where these methods are effective.